OpenStudy (anonymous):

find the exact simplified numerical valie. answer with no negative exponents, no complex fractions and either rationalize denominator or numerator where applicable sinh(ln√2)

OpenStudy (anonymous):

I got a warning that I just give you the answer, hehehe... so, I am better shut up. good luck

OpenStudy (anonymous):

do you mean \[\sin \left( \ln \sqrt{2} \right)\]?

OpenStudy (anonymous):

no sinh x (hyperbolic sine) (ln sqrt2)

OpenStudy (anonymous):

@SWAG please, help him. I don't know how to guide him without give out the answer.

OpenStudy (swag):

I am not good at math. Also try to refrain from tagging mods.

OpenStudy (anonymous):

im stuck and im not sure if im heading in the right direction but would i use (e^x - e^-x)/2?

OpenStudy (anonymous):

some mods warned me not give out the answer. If not, I will get ban. Ha!! to this kind of problem, just one line to get the answer, I have no way to obey the rule of "NOT give out the answer"@SWAG

OpenStudy (anonymous):

@SWAG

OpenStudy (anonymous):

and ln sqrt would be 1/sqrt2 ??

OpenStudy (anonymous):